Hair combs were selling at a special price of 2 for $2.50. How many combs could you purchase for $12.25
Answer by jim_thompson5910(35256)

Hair combs were selling at a special price of 2 for $2.50. How many combs could you purchase for $12.25


2/2.5 = x/12.25

2*12.25 = 2.5*x

24.50 = 2.5x

2.5x = 24.50

x = 24.50/2.5

x = 9.8

Round down to get 9

So you can buy 9 combs.